10. Rye Co. purchased a machine with a five-year estimated useful life and no salvage value for $900,000 on January 1, 1992. In its income statements, what would Rye report as the depreciation expense for 1994 (year 3 of the asset) using the double-declining-balance method?

Round up number on the nearest dollar and enter as a whole number.

Thank you!!
Answer #1
Ans. Double declining balance method:
Double declining balance depreciation rate = 2 * 1 / life of assets
2 * 1 / 5
0.40
*Depreciation = Remaining value at the beginning of the year * Double declining balance depreciation rate
Year Reamining value at the beginning (a) Depreciation (b = a*0.40) Net book value (a - b)
1 $900,000 $360,000 $540,000
2 $540,000 $216,000 $324,000
3 $324,000 $129,600 $194,400
*Net book value = a - b
*Depreciation =   Remaining value at the beginning - Residual value
*Net book value of previous year is beginning value for current year.
So the depreciation expenses for Year 3 is $129,600.